OpenStudy (anonymous):

Solve for r:

OpenStudy (agent0smith):

We need something more than that.

OpenStudy (anonymous):

Yeah, sorry, computer malfunction...

OpenStudy (anonymous):

\[V=\frac{ 1 }{ 3 } \pi h^2(3r-h)\]

OpenStudy (anonymous):

v=(1/3)pi h^2 (3r-h) 3v=pi h^2 (3r-h) 3v/pi(h^2)=3r-h [3v/pi(h^2)]+h=3r {[3v/(pi)(h^2)]+h} /3=r